OpenStudy (anonymous):

Find the x-intercept and y-intercept of 9x-5y=-9

OpenStudy (anonymous):

x-intercept means y =0 put y=o 9x=-9 so, x-intercept is x=-1 see it?

OpenStudy (anonymous):

y-intercept means x=0 if x=0 -5y=-9 y= -9/ (-5) y=9/5 this is your y-intercept. Cool?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (anonymous):

Welcome :)

OpenStudy (annon):

5y= 9x+9 y=(9/5)x + 9/5 y=mx + b b=9/5

OpenStudy (anonymous):

Thank you!

OpenStudy (annon):

:)